A passenger aircraft with a mass of 78,000 kg flies at a cruising altitude of 11,300 m at a speed of 864 km/h. How big is it


a) kinetic energy

b) potential energy in relation to the ground

c) mechanical energy?


1
Expert's answer
2021-01-16T17:22:23-0500

a) We can find the kinetic energy from the formula:


"KE=\dfrac{1}{2}mv^2,"

"KE=\dfrac{1}{2}\cdot78000\ kg\cdot(864\ \dfrac{km}{h}\cdot\dfrac{1000\ m}{1\ km}\cdot\dfrac{3600\ s}{1\ h})^2=2.25\cdot10^9\ J."

b) We can find the potential energy in relation to the ground from the formula:


"PE=mgh,""PE=78000\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot11300\ m=8.64\cdot10^9\ J."

c) By the definition of the mechanical energy, we have:


"E=PE+KE,""E=8.64\cdot10^9\ J+2.25\cdot10^9\ J=1.09\cdot10^{10}\ J."

Answer:

a) "KE=2.25\cdot10^9\ J."

b) "PE=8.64\cdot10^9\ J."

c) "E=1.09\cdot10^{10}\ J."